OpenStudy (anonymous):

Would 7/8 + 9/8 = 1 ?

OpenStudy (yamyam70):

we have the same denominator so we add the numerators. \[\frac{ 7 }{ 8 } + \frac{ 9 }{ 8 } = \frac{ 16 }{ 8 }\] Then we put the fraction in lowest terms , 8 goes in to 16 twice so ., \[\frac{ 16 }{ 8 } = \frac{ 2 }{ 1 } = 2 \]

OpenStudy (anonymous):

Oh, I get it now. Thanks.